Update on the RCT– UPS installed.

The new UPS has finally been fully installed, and is operating. After checkout tonight, the telescope is workable with two important exceptions.

The mirror compressor is not working, and the mirror appears to be resting on its supports. Brevin installed a new capacitor in the unit on saturday, so it is surprising that the compressor is failing again so soon.

The dewar is not cooling as it should, likely due to the lack of vacuum in the dewar. This is expected since the polycold unit has been off for some time.

I’ve sent requests to KPNO and to Bruce Cook to have evacuate the dewar and look at the compressor, which I hope one or the other can get to sometime this week. The telescope scheduler is off until these can be resolved.

2 thoughts on “Update on the RCT– UPS installed.

  1. strolger Post author

    The dewar has been pumped down, and the polycold is cooling properly. The mirror compressor was restarted only to have it stop a few hours later. We will look into a replacement.

  2. strolger Post author

    The scheduler is back on, and the telescope is observing. Take a look at data from tonight to determine if your program is useful with the distorted psf. If not, please set the count on your requests to zero.

    It may take a few days to repair or begin replacing the mirror compressor.

Leave a Reply

Your email address will not be published.

You may use these HTML tags and attributes: <a href="" title=""> <abbr title=""> <acronym title=""> <b> <blockquote cite=""> <cite> <code> <del datetime=""> <em> <i> <q cite=""> <strike> <strong>